OpenStudy (anonymous):

Which value for x makes the sentence true? 4x - 2 = 22 x = 28 x = 24 x = 4 x = 6

OpenStudy (black_ninja123):

4*6=24 24-2=22 x=6

OpenStudy (black_ninja123):

Same thing as the last question

OpenStudy (anonymous):

thanks again :P

OpenStudy (black_ninja123):

Your welcome

OpenStudy (lizz123):

4x - 2 = 22 4x + 2 = + 2 4x __________ = _______________ 4x = 24 __________ _____________ 4 4 x=6

OpenStudy (anonymous):

O_o woW